### v2

# Constraint Learning for Parametric Point Cloud

Abstract
Parametric point clouds are sampled from CAD shapes, and have become increasingly prevalent in industrial manufacturing. However, most existing point cloud learning methods focus on the geometric features, such as developing efficient convolution operations, overlooking the important attribute of constraints inherent in CAD shapes, which limits these methods’ ability to comprehend CAD shapes fully. To address this issue, we analyzed the effect of constraints, and proposed its deep learning-friendly representation, after that, the Constraint Feature Learning Network (CstNet) was developed to extract and leverage constraints. Our CstNet includes two stages. Stage 1 extracts constraints from B-Rep data or point cloud. Stage 2 leverages coordinates and constraints to enhance the comprehension of CAD shapes. Additionally, we built up the Parametric 20,000 Multi-modal Dataset for the scarcity of labeled B-Rep datasets. Experiments demonstrate that our CstNet achieved state-of-the-art performance on both public and proposed CAD shape datasets. To the best of our knowledge, CstNet is the first constraint-based learning method tailored for CAD shape analysis.
Project page: https://cstnetwork.github.io/
1 Introduction
Parametric point clouds are derived from parametric templates, which consist of primitives, constraints, and dimensional parameters. By specifying all parameters, the templates can be instantiated into CAD shapes, as illustrated in Fig. 1. Parametric point clouds provide detailed information that is essential for various engineering applications, such as model design and product machining.
In recent years, deep learning methods have emerged as powerful tools for point cloud analysis, with models such as PointNet [32], PointNet++ [33], and PointCNN [24] demonstrating significant success across various tasks. These models are primarily designed and evaluated on graphic objects, such as animals, trees, and other free-form shapes, where shape comprehension relies on geometric features visible in contours. However, some CAD shapes, despite exhibiting minimal visual differences, serve distinct purposes due to variations in their functional regions. This similarity in appearance poses challenges in distinguishing them based on geometric features. Notably, variations in the functional regions often introduce differing constraints, as these regions need to fit closely with other parts to perform functions such as transmission and positioning. Thus, distinguishing them from a constraint-based perspective becomes more feasible, as illustrated in Fig. 1.
Figure 1: Motivation. Some similar CAD shapes serve different functions. The eccentric wheel has a rotation axis that is offset from its outer cylindrical surface, facilitating precise positioning or tension adjustment. In contrast, the flywheel has a coaxial rotation axis with its outer cylindrical surface and is used to store rotational kinetic energy. Although these shapes are challenging to distinguish visually, they can be easily differentiated from the perspective of constraints.
Additionally, the limited availability of B-Rep datasets poses barriers to deep learning on CAD shapes. Most existing CAD shape datasets consist of mesh files, such as MCB [16] and ESB [12]. While mesh files could approximate the appearance of CAD shapes, lack crucial boundary information. In contrast, B-Rep data [19, 56] serves as the native representation of CAD shapes and is therefore more suitable for dataset construction. However, labeled B-Rep datasets remain relatively scarce, for example, FabWave [1] includes only 2,133 B-Rep files. Although the ABC [18] contains a large number of B-Rep files, it remains unlabeled.
To facilitate the constraint feature learning of CAD shapes, we conducted studies on deep learning methods and dataset development. Traditional constraint definitions are not well-suited for deep learning, as they vary across constraint types and involve multiple possible combinations to represent primitives’ relations. To address this challenge, we proposed a novel constraint representation as point-wise local attributes. Afterward, we developed the CstNet for constraint extraction and feature learning, which consist of two stages. Stage 1 is built for constraint acquisition, if B-Rep data is available, the openCASCADE Technology (OCCT) is used to extract constraints, otherwise, the constraint prediction from point cloud (CST-PCD) module is adopted. The CST-PCD is designed to rely on local features only, enabling it to generalize to unseen datasets after pre-training. Stage 2 leverages both point cloud and constraints for constraint feature learning, which employs both point-level and feature-level attention for effective feature extraction, enhancing the comprehension of CAD shapes. Finally, we built up a multi-modal classification dataset Parametric 20000, which contains B-Rep, Mesh, and point cloud. Given the current scarcity of labeled CAD shape datasets, Parametric 20000 offers researchers an expanded resource for their work. Our contributions are threefold:
  • We introduced the deep learning-friendly expression of constraints.
  • We designed the CstNet for constraint extraction and constraint feature learning.
  • We collected and curated the Parametric 20000 dataset, which consists of mesh, point cloud, and B-Rep data.
Our CstNet has been validated across various tasks, on the mechanical classification dataset MCB [16], our model achieved an overall accuracy improvement of 2.98 % compared to the state-of-the-art, and for mechanical segmentation dataset 360 Gallery [19], that is 6.74 %.
2 Related Work
For the published deep learning methods focusing on CAD shapes are very few, and even fewer aimed at parametric point clouds, this section will provide an overview of existing point cloud learning methods as well as CAD shapes-oriented models. By integrating insights from the above areas, a comprehensive review of the relevant works could be presented.
2.1 Deep learning on point clouds
Point cloud deep learning can be categorized into three types based on data processing approaches: view-based, voxel-based, and point-based methods.
View-based methods typically project the point cloud from multiple views, the obtained figures are then fed in convolutional neural networks (CNNs) for feature extraction [37, 15, 6, 54, 53, 42]. However, these methods face significant challenges when applied to large-scale scanned data, primarily due to the difficulty of covering entire scenes from single-point perspectives, and the choice of projection views significantly impacts recognition performance. Besides that, it cannot be extended to 3D data segmentation or reconstruction easily.
Voxelization-based methods convert point clouds into 3D voxels, and then employ 3D CNNs for feature extraction [30, 47, 35, 40, 17, 20]. However, the size and orientation of the voxels can affect computational efficiency and memory usage. The voxel grid quantization process may lead to geometric detail loss, and the inherent sparsity of point clouds may result in many useless voxels.
Point-based methods directly use points as input, thereby avoiding information loss associated with view-based and voxel-based methods, it is also the most popular point cloud learning method [32, 33, 24, 10, 52, 2, 38, 27, 58, 59, 9, 7, 50, 49, 48, 34, 28, 26, 41, 46, 4, 57]. Typically, symmetric functions and k-nearest neighbors (KNN) are employed for feature extraction. However, the unordered nature and transformation invariance of point clouds pose challenges for efficient feature extraction.
2.2 Deep learning on CAD shapes
CAD shapes deep learning can be categorized into three types based on the input data format: B-Rep-based learning, mesh-based learning, and sketch-based learning.
B-Rep-based methods typically involve B-Rep data conversion and feature extraction. Since B-Rep data cannot be directly fed into deep learning models, the challenge lies in converting B-Rep into deep learning compatible format, with graph representation commonly used. These approaches are commonly applied to classification [13], segmentation [19, 5], assembly [14, 43], engineering sketch generation [11], and CAD operation sequence prediction [45, 60].
Mesh-based learning methods use triangle models as input, such as STL and OBJ. Compared to B-Rep, mesh representations lack boundary details, which prevents these methods from leveraging data conversion for improved performance. As a result, contributions often rely on mechanical datasets provided by the authors. These methods are primarily used for classification [29], retrieval [3], segmentation [39], and B-Rep generation [8].
Sketch-based learning takes hand-drawn sketches as input, with outputs typically being engineering sketches [8, 44, 31, 36, 25], 3D models [21, 22, 55], or CAD operation sequences [51]. Most of these models are generative, and frequently utilize transformer modules. A key challenge is incorporating the regularity and connectivity of primitives, which are typically overlooked in graphic object analysis.
While there has been substantial progress in above domains, the engineering applications of parametric point cloud analysis demand further innovation to ensure accuracy and robustness.
3 Method
3.1 Problem Statement
Constraints are crucial for CAD shapes, can they be utilized to enhance deep learning methods’ performance? To answer above question, we designed the following validation experiments. The objective is to compare the prisms shown in Fig. 2 with cuboids. In each experiment, only prisms with a specific angle are selected, creating a binary classification task between prisms and cuboids. Consequently, eight independent experiments are conducted to cover all angles. As the prism’s angle approaches 90°90°90  °, its resemblance to cuboids increases, making it progressively more challenging to distinguish them. The constraint-aware model is built on PointNet++ backbone, this model predicts constraints and then adopts it for classification. The constraint representation is defined as the relation between the point attached primitive and a reference plane, more details in supplementary.
Figure 2: Prisms compared to cuboid.
The experimental results are presented in Fig. 3. For prisms with angle of 50°50°50  ° and 89°89°89  °, the constraints have little improvement on the results. The reason lies in the 50°50°50  ° prism differs substantially from the cuboid, while the 89°89°89  ° prism is too similar to the cuboid. For other circumstances, incorporating constraints improves classification accuracy significantly, demonstrating that the constraint enhanced the model’s comprehension of CAD shapes.
Figure 3: Classification accuracy of validation experiments.
However, the constraint learning method designed in the validation experiments is only applicable to simple scenarios. To develop a more generalized approach, we further investigated the constraint representation and deep learning architecture tailored for parametric point clouds.
3.2 Constraint Representation
Constraint should be defined as point-wise attribute for point cloud, analogous to normal vector, as point cloud is inherently unordered, assigning constraints at the point-wise level ensures that their representation is independent of point permutation.
Following the constraint definitions in CAD software is not applicable. However, once the constraint is applied, the positional relations between primitives become fixed. Therefore, if the type of primitives and their relations are determined, the constraints on CAD shapes can be defined comprehensively.
The point-wise primitive type can be represented as the type of the primitive to which the point is attached, using one-hot encoding.
The positional relation is defined as the combination of main axis direction and adjacency. Considering that the direction of primitives’ main axis is unique, as shown in Fig. 4, relations such as parallel and vertical could be derived from the direction of their main axes. However, some dimensional relations can not, such as the cylinders’ radius equivalence. Thus, additional data is required to supplement this information.
Figure 4: Main axis of primitives. Plane: normal vector. Cylinder, cone, and sphere: rotation axis.
Since the connected primitives are the basis of CAD shapes, the adjacency and main axis direction impose limitations on the distance and size of primitives. For example, in a cuboid, the adjacency and normal of the six faces ensure that the distance between the top and bottom faces equals the height of the front face, and the size of the back face is the same as the front one. Therefore, if the adjacency is combined with main axis directions, the positional relation can be determined comprehensively.
Adjacency represents the connection between primitives and cannot be expressed as point-wise attribute directly. The edges are key indicators of adjacency, with each edge corresponding to two adjacent primitives, adjacency can therefore be inferred from edges. For point cloud, an edge can be represented by marking points near it, thus, point-wise adjacency is expressed as whether the point is near an edge, which can be represented by two-channel one-hot vector.
In summary, each composition of constraints in parametric point cloud is illustrated in Fig. 5, and each point with constraint is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency).
Figure 5: Constraint representation of parametric point cloud.
3.3 Method Overview
Our CstNet and module details are depicted in the top and bottom sides of Fig. 6 respectively. It should be noted that the CST-PCD module requires pre-training, and once the pre-training is completed, its parameters are kept frozen. Given the locality of the constraints expression we designed, the pre-trained CST-PCD module is able to generalize to other unseen CAD shape datasets directly. In this work, the dataset for pre-training CST-PCD module comes from ABC [18].
Figure 6: Overview of CstNet. The top side illustrates the overall architecture of CstNet, comprising two stages. Stage 1 is designed for constraint acquisition. When B-Rep data is available, the CST-BRep module is utilized to extract constraints; otherwise, use the CST-PCD module. Stage 2 performs constraint feature learning, facilitating a deeper understanding of CAD shapes. The bottom side presents the details of module design.
3.4 Constraint Acquisition
When B-Rep data is available, the CST-BRep module can be used to compute constraints.
For the input B-Rep data, tessellation is performed by OCCT to generate a mesh model, followed by Poisson disk sampling [23] via pymeshlab to ensure a uniform point distribution and control the point count, if point cloud generation is not needed, this step could be skipped.
For adjacency computation, we identify valid edges first. In this study, an edge is considered invalid if the two adjacent primitives are smoothly connected along the edge, this work is accomplished by OCCT. Next, we compute the distance from each point to all valid edges. A point is marked as near an edge if the minimum distance is below a threshold (determined by the model surface area, set to $0.08\times Area/4\pi$ in this paper). For the primitive type and main axis direction of a point, we first determine which primitive it attached, OCCT is then used to analyze the primitive, providing both the primitive type and main axis direction.
In the absence of B-Rep data, such as working with mesh or voxel models, it can be converted into point cloud, after which the CST-PCD module can be used to predict constraints. Since each point’s constraints are local information, the computation only requires data from its neighboring points, meaning that global information is unnecessary. The constraint computation of a point $\boldsymbol{p}$ could be divided into the following steps:
Step 1. Identify neighbor points around $\boldsymbol{p}$.
Step 2. From all neighbor points, identify which belongs to the same primitive as $\boldsymbol{p}$, these points are referred to as valid points.
Step 3. Fit shapes such as cylinders or planes using valid points, and determine the primitive type with the smallest fitting error.
Step 4. Based on the valid points and the primitive type, calculate the main axis direction and assess whether the point is near an edge.
While the above process can be carried out using traditional methods, it is highly complex, which is why a deep learning model is employed to achieve the same.
The design of CST-PCD module is inspired by the aforementioned calculation process. For Step 1, which involves finding neighbor points, the traditional approach is to use the K Nearest Neighbors (KNN) algorithm, which identifies neighbors within a sphere. However, for constraint computation, it is more effective to search for neighbors along the shape’s surface. This approach captures more relevant points while reducing the influence of redundant points, as shown in Fig. 7. To this end, we proposed a neighbor searching method called SurfaceKNN, details in supplementary.
Figure 7: Comparison of KNN and SurfaceKNN. Blue point: center, red point: neighbor.
For Step 2, which involves identifying valid points from neighbors, the attention mechanism between points is employed to approximate this process, which applies different weights to neighbor points when updating the center point’s embedding. We designed the point attention layer based on vector self-attention, the relation between points is designed as embedding subtraction, as shown in Eq. 1.
$\boldsymbol{f}^{\prime}_{i}=\sum_{f_{j}\in\mathcal{N}_{i}}\rho\Bigl{(}mlp\bigl{(}\mathrm{Q}(\boldsymbol{f}_{i})-\mathrm{K}(\boldsymbol{f}_{j})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{i}).$ (1)
Where $\mathcal{N}_{i}$ is the feature collection of all neighbor points around $\boldsymbol{p}_{i}$, $\rho$ is the normalization function (softmax in this paper), and $\odot$ denotes element-wise multiplication. $\mathrm{Q}(\boldsymbol{f})$, $\mathrm{K}(\boldsymbol{f})$, and $\mathrm{V}(\boldsymbol{f})$ are defined in the same way by concatenating $\boldsymbol{f}$ with the positional encoding $\Delta$ and then feeding it into MLP, where $\Delta$ is defined as $\Delta=\boldsymbol{p}_{i}-\boldsymbol{p}_{j}$.
Step 3 and Step 4 involve determining the Primitive Type and Main Axis Direction, both steps can be accomplished by MLPs. We employed a U-Net-like structure to enhance information utilization. In the CST-PCD module, C-MLP refers to concatenating the input tensors and feeding them into MLPs. The final MLP in this module consists of three parallel MLPs with the same input, where the outputs correspond to Primitive Type, Main Axis Direction, and Adjacency.
The structure of CST-PCD module reveals that only local information is utilized, which aligns with the inherently local nature of the constraint representation we designed. For no global information is incorporated and most CAD shapes consist of a limited set of local geometry, such as plane, cylinder, and sphere, the CST-PCD can generalize to unseen datasets if pre-trained on a large dataset. As a result, our CST-PCD enables the effective use of existing unlabeled B-Rep datasets, making it possible to serve meaningful tasks.
3.5 Constraint Feature Learning
Stage 2 of CstNet is designed to leverage constraints to deepen the model’s understanding of CAD shapes. The designed architecture extracts features from Primitive Type, Main Axis Direction, and Adjacency separately, attention between features is then applied to update constraint feature. The structure and module details are illustrated in Fig. 6.
The Triple MLP takes the point cloud, Primitive Type, Main Axis Direction, and Adjacency as inputs. In this module, the point cloud is concatenated with the Primitive Type, Main Axis Direction, and Adjacency separately before being fed into three individual MLPs. The outputs are the Primitive Feature, Axis Feature, and Adjacency Feature, respectively.
In the C-MLP layer, point cloud is concatenated with Primitive Type, Main Axis Direction, Adjacency, and then feeds into MLP, with the output being the Parametric Feature.
[…]
The Fea Attention layer also takes the Primitive Feature, Axis Feature, Adjacency Feature, and Parametric Feature as input, attention between features is employed to update the constraint feature efficiently. This design considers that different points may focus on different features, for example, points near edge may focus more on the Adjacency Feature, while others may pay more attention to the Primitive or Axis Feature. The Fea Attention is developed by vector attention as formulated in Eq. 2.
$\boldsymbol{f}^{\prime}_{cst}=\sum_{\boldsymbol{f}_{j}\in\mathcal{F}}\rho\Bigl{(}mlp\bigl{(}\mathrm{Q}(\boldsymbol{f}_{cst})-\mathrm{K}(\boldsymbol{f}_{j})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{cst}).$ (2)
In which =$\mathcal{F}=\{\boldsymbol{f}_{pmt},\boldsymbol{f}_{axs},\boldsymbol{f}_{adj},\boldsymbol{f}_{cst}\}$, $\boldsymbol{f}_{pmt}$ is Primitive Feature, $\boldsymbol{f}_{axs}$ is Axis Feature, $\boldsymbol{f}_{adj}$ is Adjacency Feature, and $\boldsymbol{f}_{cst}$ is Constraint Feature. $\Delta$ is defined as one-hot encoding of $\boldsymbol{f}_{j}$ type.
For downstream tasks, if classification is required, the final constraint feature can be fed into MLP and softmax. For segmentation tasks, upsampling and skip link concatenation similar to those in PointNet++ [33] can be applied.
4 Experiments
4.1 Experimental Settings
In general, we train CstNet using the Negative Log Likelihood Loss (NLLLoss) and optimized by Adam optimization for 200 epochs, with a weight decay of 0.0001 and an initial learning rate of 0.0001. We use StepLR learning rate scheduler, with step size of 20 and gamma of 0.7. The batch size is 16. All experiments are performed on GeForce RTX 4090 GPU.
Parametric 20000 dataset: our Parametric 20000 dataset contains 19,739 multi-modal instances categorized into 75 classes. Each instance includes B-Rep data, mesh, and point cloud. All B-Rep data are collected from the following three sources. More details in the supplementary.
Sources 1 (36.05 %): Downloaded and purified from TraceParts [16]. The CAD shapes on TraceParts are collected from manufacturing companies.
Sources 2 (24.12 %): Provided by a CAD company [^1] https://www.newdimchina.com/ that collaborates with our team. These shapes are all mechanical components.
Sources 3 (39.83 %): Instantiated from parametric templates designed by our team. These templates are designed based on common mechanical structures.
Our Parametric 20000 dataset encompasses various common formats of CAD shapes, making it suitable for a wide range of tasks. The careful selection and purification process undertaken by our team further ensures its high quality.
4.2 Constraint Acquisition
The extraction of constraints from B-Rep data is accomplished by CST-BRep module, with results visualized in the "GT" column of Fig. 8. By referring to the model row, this module performs well in extracting three components of constraint.
Extracting constraints from point cloud is achieved by CST-PCD module. The training set is sourced from 25 trunks of ABC [18], with each trunk containing 10,000 B-Rep files. We utilized CST-BRep module to extract constraints as Ground Truth for model training. After the training finished, the prediction accuracy of Primitive Type reached 94.23 %, for Adjacency, that is 94.62 %, and the MSE loss for Main Axis Direction is 0.0559.
In subsequent classification and segmentation tasks, the above pre-trained CST-PCD module is utilized to predict constraints with its weights frozen and no further updates. To ensure that only point cloud is used as input across all methods in comparison, even when B-Rep data is available, the CST-BRep module will not be activated to obtain constraints.
After CST-PCD module was pre-trained, we evaluated it on a new trunk of ABC. The visualization of constraint prediction is shown in the "Pred." column of Fig. 8. After traversing this trunk, the prediction accuracy of Primitive Type reached 91.19 %, for Adjacency, that is 90.24 %, the MSE loss of Main Axis Direction is 0.0998, above demonstrating the strong capability for constraint prediction.
Figure 8: Constraint prediction on ABC. MAD: Main Axis Direction. PT: Primitive Type. ADJ: Adjacency (whether point is near edge). Main Axis Direction has been processed for visualization, details in supplementary.
Additionally, we evaluated the pre-trained CST-PCD module on the CAD shape dataset MCB [16] to predict constraints. For the MCB consists of mesh files, it was not possible to extract Ground Truth constraints for comparison. Based on the results visualized in Fig. 9, CST-PCD demonstrated a high level of constraint prediction accuracy even on an unseen dataset.
Figure 9: Constraint prediction on MCB.
4.3 Classification
The classification results on the Parametric 20000 dataset are presented in Tab. 1, where our CstNet outperformed other methods across both metrics. Additionally, the instance accuracy of all methods is below 90 %, with a notable performance gap between them, indicating that this dataset poses challenges and effectively differentiates between methods.
Table 1: Classification results on Parametric 20000. Acc: accuracy over instance %, Acc*: accuracy over class %, F1: F1-score, mAP: mean average precision %.
Method Acc Acc* F1 mAP
PointCNN [24] 61.92 55.29 53.97 40.99
PointNet [32] 75.61 74.08 74.23 82.15
PointNet++ [33] 82.89 86.80 82.31 88.05
PointConv [46] 80.53 82.10 81.86 86.54
DGCNN [41] 85.99 88.20 87.20 89.20
3DGCN [26] 86.19 86.31 85.37 87.64
Ours 89.93 91.04 90.32 92.32
We conducted evaluations on the CAD shape dataset MCB [16], with results presented in Tab. 2. Our CstNet achieved the highest scores across all metrics on MCB-A, but performed less effectively on MCB-B. This discrepancy may stem from the composition of MCB: MCB-A includes shapes from TraceParts (69.27 %), 3D Warehouse (21.70 %), and GrabCAD (9.03 %), while MCB-B is limited to shapes from 3D Warehouse and GrabCAD. TraceParts provides data from manufacturing companies, whereas 3D Warehouse and GrabCAD contain shapes for general usage [16]. As the CST-PCD module in CstNet is pre-trained on ABC, which primarily comprises mechanical components, this pre-training may lead to poor constraint prediction performance on MCB-B, thereby influencing the classification results.
Table 2: Classification results on MCB. A / B column: results of MCB A / B subset.
Method Acc.instance % Acc.class % F1-score mAP %
A B A B A B A B
PointCNN [24] 93.89 93.67 81.85 86.80 83.86 88.63 90.13 93.86
PointNet [32] 86.78 80.36 67.70 65.30 86.55 79.89 74.08 71.66
PointNet++ [33] 87.45 93.91 73.68 87.97 88.32 73.45 91.33 91.37
SpiderCNN [52] 93.59 89.31 79.70 79.29 81.30 80.72 86.64 82.47
PointConv [46] 93.25 90.64 80.24 80.17 71.31 76.07 82.19 86.09
DGCNN [41] 92.54 90.75 74.47 78.17 76.12 79.69 74.27 79.73
3DGCN [26] 93.71 92.13 78.71 84.31 84.59 79.24 84.35 84.35
Ours 96.87 92.61 89.21 85.94 89.85 85.68 93.17 91.25
4.4 Segmentation
We conducted experiments on the mechanical segmentation dataset 360 Gallery [19], which contains point clouds with segmentation labels based on CAD modeling operations. However, this labeling approach presents disadvantages for point cloud segmentation. For example, a cube in this dataset has four faces labeled as "ExtrudeSide" and two faces labeled as "ExtrudeEnd", which poses an unrealistic task for point cloud networks. Additionally, the distribution of points across different labels is highly imbalanced. For instance, points labeled as "ExtrudeSide" account for 45.94 %, while those labeled as "RevolveEnd" make up only 0.12 %. Despite these limitations, to our knowledge, 360 Gallery is currently the only publicly available point cloud dataset for CAD shape segmentation tasks.
The segmentation results are presented in Tab. 3, where our method outperforms the state-of-the-art methods on most metrics. The results also highlight the limitations of 360 Gallery in terms of point cloud segmentation. Apart from "ExtrudeSide" and "ExtrudeEnd", none of the models were able to differentiate the other segmentation categories except DGCNN managed to predict "RevolveSide" with a low mIoU.
Table 3: Part Segmentation Results on 360 Gallery. PN: PointNet. PN2: PointNet++. DGN: DGCNN. 3DN: 3DGCN. OA: overall accuracy %. mIoU: instance mIoU %. mIoU*: class mIoU %. ESide: Extrude Side mIoU %, EEnd: Extrude End mIoU %. Other segmentation categories’ mIoU are all 0 except RevolveSide mIoU of DGCNN is 9.60 %, more datails in supplementary.
Method OA mIoU mIoU* ESide EEnd
PN [32] 42.13 18.89 7.04 19.83 36.47
PN2 [33] 58.18 26.98 11.27 49.80 40.36
DGN [41] 64.02 35.54 14.36 54.48 50.82
3DN [26] 60.58 33.03 12.23 50.67 47.20
Ours 70.76 46.65 20.51 61.29 60.65
A portion of the segmentation results on 360 Gallery test set are shown in Tab. 4, where our CstNet shows segmentation results closest to the Ground Truth. PointNet’s segmentation differs from other methods, for instance, in the third and fourth rows, PointNet mistakenly segments a single primitive into two different categories. This discrepancy likely arises from the model’s inability to effectively capture local features.
Table 4: Visualization of part segmentation on 360 Gallery. GT: ground truth. PN: PointNet [32]. PN2: PointNet++ [33]. DGN: DGCNN [41]. 3DN: 3DGCN [26].
GT PN PN2 DGN 3DN Ours
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
4.5 Ablation Studies
Our model leverages all three components of constraint: Primitive Type, Main Axis Direction, and Adjacency. Additionally, we utilized the SurfaceKNN to obtain more suitable neighbors for each point. To validate the effectiveness of our design, we conducted a series of experiments on Parametric 20000 dataset.
For each component of constraint, we evaluated their importance by disabling one at a time. This operation is performed by setting it as a zero tensor before feeding into the model. The results are presented in Tab. 5, which show that all three components contribute to improving the performance of our CstNet.
To assess the effectiveness of SurfaceKNN, we compared it with KNN. From Tab. 5, it can be concluded that in most cases, our SurfaceKNN outperforms KNN. However, when the Primitive Type is disabled, the results are similar between the two methods, indicating that SurfaceKNN primarily enhances the effectiveness of the Primitive Feature.
Table 5: Effects on Parametric 20000 dateset with varying neighbor search and constraints. Row 3 - 6: SurfaceKNN, Row 7 - 10: KNN.
Constraint Acc Acc* F1 mAP
MAD ADJ PT
89.93 91.04 90.32 92.32
86.32 89.05 88.38 90.59
88.14 89.61 88.64 91.35
88.68 88.94 89.01 91.22
89.03 90.00 89.22 91.43
85.04 85.93 85.74 88.08
87.27 88.23 87.98 90.00
88.73 88.97 88.91 91.21
5 Conclusions
In this paper, we studied the CAD shapes and introduced the constraints expression which is suitable for deep learning methods. Afterward, we proposed the CstNet for extracting and learning constraint features. Additionally, we built up a multimodal CAD shape classification dataset Parametric 20000. Experimental results on both the public datasets and proposed dataset demonstrate that our CstNet achieves state-of-the-art performance across multiple metrics. Comprehensive ablation studies further validate that each component of constraint and SurfaceKNN contributes to improve the performance of CstNet.
CAD shapes possess numerous attributes. In this paper, we analyze the constraints from a functional perspective. In future work, we plan to investigate additional aspects such as part design and manufacturing processes, to achieve a deeper understanding of CAD shapes.